Use a contour to evaluate S (2-1)723 +1) da 7T Answer:
Answer #1

# Am! I= 8 9 20249 da 736-1)(3341) 0 2 ²422 let fr) = C-1)(734) poles of free where as +1+113 9 given by: (2-1)(4340)20 +7=1Since a -R 1 -0-1 1 48R 0 -R 471 71 + IR . faida R 121² 2121 1d2l C = (-R,=>JU CIUTAT, I-81) UC 01148!,p] UCH Hence by Resur